Suppose the simplified consolidated balance sheet shown below is for the entire commercial banking system and that all figureWhat is the maximum amount the banking system might lend? Answer is complete but not entirely correct 1,000 billion Show in cPlease help me complete the chart as well! thank you
Answer #1

Ans:

A). I). The amount of excess reserve in this commercial building systems=

Excess reserve= total reserve- required reserve

Required reserve= reserve ratio x checkable deposits

= 10% * 200 billions

= 20 billion

Excess reserve= 58 billion - 20 billion

= 38 billion

II). Maximum amount that banking system will lend = 1/ reserve ratio* excess reserve

=1/10%* 38 billion

= $380 billion

III). Money multiplier= 1/ reserve ratio

= 1/ 0.10

= 10

B).

I). The amount of excess reserve in this commercial building systems=

Excess reserve= total reserve- required reserve

Required reserve= reserve ratio x checkable deposits

= 5% * 200 billions

= 10 billion

Excess reserve= 58 billion - 10 billion

= 48 billion

II). Maximum amount that banking system will lend = 1/ reserve ratio* excess reserve

=1/10%* 48 billion

= $480 billion

III). Money multiplier= 1/ reserve ratio

= 1/ 0.05

= 20

The difference between the amount of lending is ($480 billion -380 billion)

= $100 billion

Chart:

Assets Liabilities
Particulars 1 A). B). Particulars 1. A). B)
Reserves 58 56 45 Checkable deposits 200 580 680
Securities 42 42 42
Loans 100 380 480